With all the fad diets advertised on TV, a researcher wanted to know if there was a significant difference in weight loss depending on which new fad diet you were on. Over the course of 2 weeks, a total of 12 male volunteers who weighed the same and had similar lifestyles were broken up into 3 groups: Diet 1, Diet 2, and Diet 3. Conduct an ANOVA on the following table with alpha = 0.5 to determine if at least one of the diets has significantly different weight loss compared to the other two.

1) What are the null and alternative hypotheses? A.) H0: Mean weight loss from Diet 1 = mean weight loss from Diet 2 = mean weight loss from Diet 3 4 HA: Mean weight loss from Diet 1 > mean weight loss from Diet 2 > mean weight loss from Diet 3 B.) H0: Mean weight loss from Diet 1 = mean weight loss from Diet 2 = mean weight loss from Diet 3 HA: At least one mean weight loss is significantly different C.) H0: At least one mean weight loss is significantly different HA: Mean weight loss from Diet 1 > mean weight loss from Diet 2 > mean weight loss from Diet 3 D.) H0: Mean weight loss from Diet 1 = mean weight loss from Diet 2 = mean weight loss from Diet 3 HA: Mean weight loss from Diet 1 < mean weight loss from Diet 2 < mean weight loss from Diet 3

2.) What are the dfbetween? a. 2 b. 3 c. 11 d. 12

3.) What are the dfwithin? a. 9 b. 10 c. 11 d. 12

4.) What is the critical value? (Round to 4 decimal places) a. 3.8853 b. 3.9823 c. 4.2565 d. 3.5874
Answer #1

QUESTION 1

Here ,option B is correct here about hypothesis problem

Question 2

df(between) = 2

Question 3

df(within) = (4-1) *3 = 9

Question 4

Critical value for dF1 = 2, dF2 = 9

F(critical) = F(0.05, 2, 9) = 4.256

NOTE : I hope you are happy with my answer

*  Please help me with your rating  *

     THANK YOU